# Ambig-SWE: Interactive Agents to Overcome Underspecificity in Software Engineering

### v2

@@ 1 Introduction
The multi-stage evaluation allows for targeted improvements in individual aspects, offering a pathway to enhance overall system performance. Through our evaluations across the different settings, we find that interactivity can boost performance on underspecified inputs by up to 74% over the non-interactive settings, though performance varies across models (§3). LLMs default to non-interactive behavior without explicit encouragement, and even with it, they struggle to distinguish between underspecified and well-specified inputs. Claude Sonnet 4 and Claude Sonnet 3.5 are the only evaluated LLMs that achieve notable accuracy (89% and 84%, respectively) in making this distinction. Prompt engineering offers limited improvement, and its effectiveness varies across models (§4). When interacting, LLMs generally pose questions capable of extracting relevant details, but some models, such as Llama 3.1 70B, fail to obtain sufficient specificity (§4). As models grow more capable (e.g., from Claude Sonnet 3.5 to Claude Sonnet 4), interaction provides diminishing returns, suggesting current training practices may not adequately leverage clarification. In summary, this study underscores the importance of interactivity in LLMs for agentic workflows, particularly in real-world tasks where prompt quality varies significantly.
@@ 3.1 Experimental setup
By default, coding agents are restricted to 30 interaction turns to produce a solution patch; however, Claude Sonnet 4 and Qwen 3 Coder are allocated up to 100 turns to account for their greater reasoning and planning capacity. In this experiment, each model is tested in the Hidden [^4] Claude Sonnet 4 is evaluated on a subset of 100/500 instances in the Hidden setting. The model compensates for the lack of information with increased exploration and solution attempts leading to substantially higher evaluation costs. The findings are still statistically significant (Table 4)., Interaction, and Full settings to evaluate its ability to leverage interaction and optimize performance on underspecified issues. The results, as shown in Figure 3, confirm the expected increase in resolve rates for all models as more information becomes available to the agent. The difference between the Hidden and Interaction settings is significant for all evaluated models (Table 4), emphasizing the impact of interaction on task completion. The performance gap between the Interaction and Full settings is also significant across all models, highlighting the unrealized potential. Specifically, for the Hidden vs. Interaction settings, proprietary models show stronger evidence of a significant difference. These results suggest that the ability to leverage interaction varies across models, with proprietary models generally demonstrating greater effectiveness in utilizing interaction compared to open-weight models. However, as open-weight models improve, they can even outperform proprietary models with interaction, as demonstrated by Qwen 3 Coder.
@@ 3.2 Leveraging interaction in underspecificity
Using interaction, the Claude Sonnet 3.5 models and Haiku 3.5 recover up to 80% of the performance in the Full setting. However, with Deepseek, and Llama 3.1, the relative performance is lower, of 59%, and 54%, respectively. Claude Sonnet 4’s relative performance (61%) is lower than that of its predecessor, and absolute performance with interaction is also similar. Some models achieve higher resolve rates in the Hidden setting likely due to their superior programming acumen, or data leakage. Better programming models can potentially extract more information from the stack trace by reproducing the error themselves. Claude Sonnet 4 extensively explores the codebase and attempts multiple solutions to overcome the lack of information in the Hidden setting. On the other hand, Qwen 3 Coder displays unique behavior in this setting and relies on its internal knowledge for key insights about missing information (example in §A.7). These correct assumptions might inflate its performance in this setting. We observe that the Claude Haiku model achieves a performance relative to the Full setting similar to that of Claude Sonnet 3.5, despite having inferior coding abilities. Thus, there does not seem to be a direct correlation between the number of parameters or coding ability and a model’s ability to leverage interaction. This hints towards better training practices that can lead to better integration of the new information.
@@ 3.3 Impact of interaction details on model performance
Claude Sonnet 4 27.75 43.75 46.55
Qwen 3 Coder 18.60 47.60 40.20
@@
Takeaway: While proprietary models like Claude Sonnet 3.5 and Haiku 3.5 effectively leverage interaction (recovering  80% of the performance gap), recent capable models show a shift: Claude Sonnet 4’s relative gains diminish despite stronger absolute performance, and Qwen 3 Coder rigidly adheres to predetermined protocols even when users provide explicit guidance. These patterns suggest that as models grow more capable, current training practices may inadequately prepare them to dynamically integrate interactive information, highlighting the need for approaches that prioritize adaptive behavior over task completion alone.
@@ 7 Conclusion, Limitations, and Future Work
  • With a brief round of clarification, leading proprietary models recover much of their fully-specified performance, while earlier open-weight models lag. Recent capable models blur this distinction. However, as models grow more capable, relative gains from interaction diminish, suggesting current training practices inadequately prepare models to dynamically integrate interactive information.
@@
Overall, a gap remains between underspecified and fully specified resolution rates. Closing it will require open-weight models to adopt stronger interaction strategies and proprietary models to engage more proactively. As models are trained to perform longer horizon tasks, they must still be trained to appropriately incorporate user inputs into the overall solution. Ambig-SWE provides a blueprint for decomposing resolution into multiple steps, enabling finer-grained analysis of where models succeed or fail. While we focus on software engineering, the methods and insights can extend to other complex, real-world agentic tasks. Thus, our work offers both a diagnostic framework for agent evaluation and a roadmap toward more robust, adaptive, and user-aligned agents that can thrive in underspecified and dynamic environments.
@@ A.3.1 Wilcoxon Signed-Rank test
Qwen 3 Coder Hidden vs Interaction 6.87e-17
Interaction vs Full 5.46e-26
Claude Sonnet 4 Hidden vs Interaction 0.03225
Interaction vs Full 9.0e-29

### v3

@@ 1 Introduction
The multi-stage evaluation allows for targeted improvements in individual aspects, offering a pathway to enhance overall system performance. Through our evaluations across the different settings, we find that interactivity can boost performance on underspecified inputs by up to 74% over the non-interactive settings, though performance varies across models (§3). LLMs default to non-interactive behavior without explicit encouragement, and even with it, they struggle to distinguish between underspecified and well-specified inputs. Claude Sonnet 4 and Claude Sonnet 3.5 are the only evaluated LLMs that achieve notable accuracy (89% and 84%, respectively) in making this distinction. Prompt engineering offers limited improvement, and its effectiveness varies across models (§4). When interacting, LLMs generally pose questions capable of extracting relevant details, but some models, such as Llama 3.1 70B, fail to obtain sufficient specificity (§4). As model capability scales (e.g., from Claude Sonnet 3.5 to Claude Sonnet 4), we observe substantial performance gains in both interactive and non-interactive settings, driven by improvements in underlying competencies that also enable more effective integration of information acquired through interaction. In summary, this study underscores the importance of interactivity in LLMs for agentic workflows, particularly in real-world tasks where prompt quality varies significantly.
@@ 3.1 Experimental setup
By default, coding agents are restricted to 30 interaction turns to produce a solution patch; however, Claude Sonnet 4 and Qwen 3 Coder are allocated up to 100 turns to account for their greater reasoning and planning capacity. In this experiment, each model is tested in the Hidden [^4] Claude Sonnet 4 is evaluated on a subset of 100/500 instances in the Hidden setting. The model compensates for the lack of information with increased exploration and solution attempts leading to substantially higher evaluation costs. The findings are still statistically significant (Table 4)., Interaction, and Full settings to evaluate its ability to leverage interaction and optimize performance on underspecified issues. The results, as shown in Figure 3, confirm the expected increase in resolve rates for all models as more information becomes available to the agent. The difference between the Hidden and Interaction settings is significant for all evaluated models (Table 4), emphasizing the impact of interaction on task completion. The performance gap between the Interaction and Full settings is also significant across all models, highlighting the unrealized potential. Specifically, for the Hidden vs. Interaction settings, proprietary models show stronger evidence of a significant difference. These results suggest that the ability to leverage interaction varies across models, with proprietary models generally demonstrating greater effectiveness in utilizing interaction compared to open-weight models. However, as open-weight models continue to advance, they narrow the gap with proprietary systems in their ability to effectively leverage interaction, as illustrated by Qwen 3 Coder.
@@ 3.2 Leveraging interaction in underspecificity
Using interaction, the Claude Sonnet 3.5 models and Haiku 3.5 recover up to 80% of the performance in the Full setting. However, with Deepseek, and Llama 3.1, the relative performance is lower, of 59%, and 54%, respectively. Claude Sonnet 4 attains the highest relative performance (89%), reflecting a stronger ability to integrate information acquired through interaction, likely supported by improved long-horizon reasoning. Efficiency, measured by action steps, provides an additional lens. Qwen 3 Coder uses roughly the same number of steps (65) in both Hidden and Interaction settings, while Claude Sonnet 4 increases its average steps from 65 to 75 when interaction is enabled. Thus, interaction improves effectiveness but not efficiency under current patterns. Some models achieve higher resolve rates in the Hidden setting likely due to their superior programming acumen, or data leakage. Better programming models can potentially extract more information from the stack trace by reproducing the error themselves. Claude Sonnet 4 extensively explores the codebase and attempts multiple solutions to overcome the lack of information in the Hidden setting. On the other hand, Qwen 3 Coder displays unique behavior in this setting and relies on its internal knowledge for key insights about missing information (example in §A.7). These correct assumptions potentially inflate its performance in this setting. We observe that the Claude Haiku model achieves a performance relative to the Full setting similar to that of Claude Sonnet 3.5, despite having inferior coding abilities. Thus, there does not seem to be a direct correlation between the number of parameters or coding ability and a model’s ability to leverage interaction. This hints towards better training practices that can lead to better integration of the new information.
@@ 3.3 Impact of interaction details on model performance
Claude Sonnet 4 12.24 60.82 67.24
Qwen 3 Coder 18.58 55.43 52.38
@@
Takeaway: Proprietary models such as Claude Sonnet 3.5 and Haiku 3.5 effectively leverage interaction, recovering up to 80% of the performance gap, and newer high-capability models (e.g., Claude Sonnet 4 and Qwen 3 Coder) further improve in combining interaction with stronger underlying competencies. However, Qwen 3 Coder often rigidly follows predetermined protocols despite explicit corrective guidance, and interaction yields no efficiency gains, as higher effectiveness is not accompanied by fewer steps. These trends suggest that, as capability scales, current training paradigms may insufficiently promote adaptive integration of interactive feedback, highlighting the need for training approaches that do not prioritize task completion alone.
@@ 7 Conclusion, Limitations, and Future Work
  • With a brief round of clarification, leading proprietary models recover much of their fully-specified performance, while earlier open-weight models lag. Recent capable models blur this distinction. However, although interaction improves task success as capability scales, it does not yield efficiency gains, indicating that current training practices insufficiently optimize for effective integration of interactive feedback.
@@
Overall, a gap remains between underspecified and fully specified resolution rates. Closing it will require open-weight models to adopt stronger interaction strategies and proprietary models to engage more proactively. As models are trained to perform longer horizon tasks, they must still be trained to appropriately incorporate user inputs into the overall solution effectively and efficiently. Ambig-SWE provides a blueprint for decomposing resolution into multiple steps, enabling finer-grained analysis of where models succeed or fail. While we focus on software engineering, the methods and insights can extend to other complex, real-world agentic tasks. Thus, our work offers both a diagnostic framework for agent evaluation and a roadmap toward more robust, adaptive, and user-aligned agents that can thrive in underspecified and dynamic environments.
@@ A.3.1 Wilcoxon Signed-Rank test
Qwen 3 Coder Hidden vs Interaction 3.54e-07
Interaction vs Full 8.91e-14
Claude Sonnet 4 Hidden vs Interaction 6.34e-18
Interaction vs Full 9.12e-06